OpenStudy (anonymous):

Need help with the substitute rule i'll post the pic up in a sec please and thank you :D

OpenStudy (anonymous):

number 47

OpenStudy (anonymous):

\[\int x(2x+5)^8~dx\] Let \(u=2x+5\), then \(\dfrac{1}{2}du=dx\). Also, solving for \(x\) in the sub, you have \(x=\dfrac{u-5}{2}\): \[\frac{1}{2}\int\left(\frac{u-5}{2}\right)u^8~du\]

OpenStudy (anonymous):

i don't get the x part

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (isaiah.feynman):

What don't you understand?

OpenStudy (anonymous):

the x part

OpenStudy (isaiah.feynman):